# Construction and classification of differential symmetry breaking operators for principal series representations of the pair $(SO_{0}(4,1),SO_{0}(3,1))$ for special parameters

### v1

@@ 7.2 Phase 3: Proving $f_{-0}=f_{+0}$
Now that we have completed Phase 2, i.e., that we have obtained $f_{-N},\ldots,f_{0}$ in the -- side, and $f_{0},\ldots,f_{N}$ in the $+$ side, the only task that remains is to check that the expressions obtained for $f_{0}$ in both sides (that we can denote by $f_{-0}$ and $f_{+0}$ respectively) coincide.
@@
Suppose $\Gamma_{0}^{-}\neq 0$ and suppose that $f_{\pm j}$ are given by (7.12) and (7.13) respectively. Then, (7.22) holds if and only if the tuple $(\widetilde{q_{N}^{-}},\widetilde{q_{N}^{+}})$ satisfies the following relation.
@@
If $a\geq 2N$, then $\deg(\widetilde{C}_{a-N}^{\lambda+N-1})=a-N>\deg(P_{0})=N-s$. Thus, the top terms of both sides of (7.24) coincide if and only if $(\widetilde{q_{N}^{-}},\widetilde{q_{N}^{+}})$ satisfies (7.23). Then, from (7.24) we have that
@@
The first case is covered by Lemma 7.9, so we obtain the same result. On the other hand, case (b) can be proved by following the same argument in the proof of Lemma (7.9). This can be done since as $s=2N-a,\ldots,N$, we have $N-s<a-N$. So in this case we obtain the same result as Lemma 7.10.
@@
$f_{\pm j}\in\text{{Pol}}_{a-N\pm j}[t]_{\text{{{even}}}},\enspace\text{ for }j=0,1,\ldots,N.$ (7.25)
@@
(1) Suppose now that $0\leq a\leq N$. Since $a-j-N<0$ for all $j=1,2,\ldots,N$, we have that $f_{-j}\equiv 0$ for all $j=1,2,\ldots,N$ from (7.25). Hence, all the functions in the -- side vanish, so it suffices to determine $f_{j}$ (for all $j=0,1,\ldots,N$). In particular, there is no Phase 3 in this case. Moreover, also by (7.25), we deduce that $f_{j}\equiv 0$ for $j=0,\ldots,N-a-1$, so the non-vanishing functions are $f_{N-a},f_{N-a+1},\ldots,f_{N}$.
@@
8 Duality Between $m$ and $-m$
In this section, we establish a duality between the F-systems for $m\geq N$ and for $m\leq-N$ (Proposition 8.2). This result, combined with Theorem 5.1, allows us to show Theorems 1.2 and 1.3 for $m=-N$. We begin by introducing some notation.
@@ 8 Duality Between $m$ and $-m$
$\begin{gathered}\psi^{+}((g_{k})_{k=p-N}^{p+N})(\zeta)=\begin{pmatrix}\displaystyle{(T_{a-p+N}g_{p-N})(\zeta)(\zeta_{1}+i\zeta_{2})^{p-N}}\\[4.0pt] \vdots\\[4.0pt] \displaystyle{(T_{a-p-N}g_{p+N})(\zeta)(\zeta_{1}+i\zeta_{2})^{p+N}}\end{pmatrix},\\[4.0pt] \psi^{-}((g_{k})_{k=p-N}^{p+N})(\zeta)=\begin{pmatrix}\displaystyle{(T_{a-p-N}g_{p+N})(\zeta)(\zeta_{1}-i\zeta_{2})^{p+N}}\\[4.0pt] \vdots\\[4.0pt] \displaystyle{(T_{a-p+N}g_{p-N})(\zeta)(\zeta_{1}-i\zeta_{2})^{p-N}}\end{pmatrix}.\end{gathered}$
@@
$\begin{gathered}\left(\widehat{d\pi_{(\sigma^{2N+1},\lambda)^{*}}}(C_{1}^{+})\otimes\text{{id}}_{{\mathbb{C}}_{p,\nu}}\right)\psi^{+},\quad\left(\widehat{d\pi_{(\sigma^{2N+1},\lambda)^{*}}}(C_{1}^{+})\otimes\text{{id}}_{{\mathbb{C}}_{-p,\nu}}\right)\Phi(\psi^{+}),\end{gathered}$
@@
$=0\quad(\text{by hypothesis}).$
@@
(1) From Proposition 8.2 above, we deduce that the space $\text{{Sol}}(\mathfrak{n}_{+},\sigma_{\lambda}^{2N+1},\tau_{-N,\nu})$ has dimension at most one and that the necessary and sufficient condition on the parameters $(\lambda,\nu,N)$ for it to be non-zero is the same as in the case $m=N$, namely $\nu-\lambda\in{\mathbb{N}}$. This establishes Theorem 1.2 for $m=-N$.
(2) As for the differential operators $\mathbb{D}_{\lambda,\nu}^{N,-N}$, from the result above, we deduce that the operator $\mathbb{D}_{\lambda,\nu}^{N,-N}$ can be obtained by reordering the scalar operators in each coordinate and replacing $i$ by $-i$ in the expression of the operator $\mathbb{D}_{\lambda,\nu}^{N,N}$. Specifically, from the expression of the map $\Phi_{\text{Sol}}^{+}$, $\mathbb{D}_{\lambda,\nu}^{N,-N}$ can be obtained by swapping the $s$-coordinate with $(-1)^{s}$ times the $(2N-s)$-coordinate and replacing $i$ by $-i$ in the operator $\mathbb{D}_{\lambda,\nu}^{N,N}$. This proves Theorem 1.3 for $m=-N$.
@@
Note that the results of this section can be applied when solving Problems A and B for $|m|>N$. In fact, it suffices to solve them when $m>N$ and then use Proposition 8.2 to obtain the solution for $m<-N$.
@@ 9.1 Renormalized Gegenbauer polynomials
$C_{\ell}^{\mu}(z)$ $:=\sum_{k=0}^{[\frac{\ell}{2}]}(-1)^{k}\frac{\Gamma(\ell-k+\mu)}{\Gamma(\mu)k!(\ell-2k)!}(2z)^{\ell-2k}.$
@@ 9.2 Algebraic properties of renormalized Gegenbauer polynomials and imaginary Gegenbauer operators
Let $\mu\in{\mathbb{C}}$ and $\ell\in{\mathbb{N}}$ and suppose that $02\mu+\ell=$ 0. Then, for any $n,d\in{\mathbb{N}}$ we have:
@@
The proof is straightforward from the definition of $S^{\mu}_{\ell}$ and the condition $02\mu+\ell=$ 0. ∎

### v2

@@ 7.2 Phase 3: Proving $f_{-0}=f_{+0}$
Now that we have completed Phase 2, i.e., that we have obtained $f_{-N},\ldots,f_{0}$ in the - side, and $f_{0},\ldots,f_{N}$ in the ++ side, the only task that remains is to check that the expressions obtained for $f_{0}$ in both sides (that we can denote by $f_{-0}$ and $f_{+0}$ respectively) coincide.
@@
[…]
@@
[…]
@@
The first case is covered by Lemma 7.9, so we obtain the same result. On the other hand, case (b) can be proved by following the same argument in the proof of Lemma (7.9). This can be done since as $s=2N-a,\ldots,N$, we have Ns<aNN-s<a-N. So in this case we obtain the same result as Lemma 7.10.
@@
$f_{\pm j}\in\text{{Pol}}_{a-N\pm j}[t]_{\text{{{even}}}},\kern 4.62497pt\text{ for }j=0,1,\ldots,N.$ (7.25)
@@
(1) Suppose now that $0\leq a\leq N$. Since ajN<0a-j-N<0 for all $j=1,2,\ldots,N$, we have that $f_{-j}\equiv 0$ for all $j=1,2,\ldots,N$ from (7.25). Hence, all the functions in the - side vanish, so it suffices to determine $f_{j}$ (for all $j=0,1,\ldots,N$). In particular, there is no Phase 3 in this case. Moreover, also by (7.25), we deduce that $f_{j}\equiv 0$ for $j=0,\ldots,N-a-1$, so the non-vanishing functions are $f_{N-a},f_{N-a+1},\ldots,f_{N}$.
@@
8 Duality Between $m$ and m-m
In this section, we establish a duality between the F-systems for $m\geq N$ and for $m\leq-N$ (Proposition 8.2). This result, combined with Theorem 5.1, allows us to show Theorems 1.2 and 1.3 for m=Nm=-N. We begin by introducing some notation.
@@ 8 Duality Between $m$ and m-m
[…]
@@
$\begin{gathered}\left(\widehat{d\pi_{(\sigma^{2N+1},\lambda)^{*}}}(C_{1}^{+})\otimes\text{{id}}_{{\mathbb{C}}_{p,\nu}}\right)\psi^{+},\qquad\left(\widehat{d\pi_{(\sigma^{2N+1},\lambda)^{*}}}(C_{1}^{+})\otimes\text{{id}}_{{\mathbb{C}}_{-p,\nu}}\right)\Phi(\psi^{+}),\end{gathered}$
@@
$=0\qquad(\text{by hypothesis}).$
@@
(1) From Proposition 8.2 above, we deduce that the space $\text{{Sol}}(\mathfrak{n}_{+},\sigma_{\lambda}^{2N+1},\tau_{-N,\nu})$ has dimension at most one and that the necessary and sufficient condition on the parameters $(\lambda,\nu,N)$ for it to be non-zero is the same as in the case $m=N$, namely $\nu-\lambda\in{\mathbb{N}}$. This establishes Theorem 1.2 for m=Nm=-N.
(2) As for the differential operators $\mathbb{D}_{\lambda,\nu}^{N,-N}$, from the result above, we deduce that the operator $\mathbb{D}_{\lambda,\nu}^{N,-N}$ can be obtained by reordering the scalar operators in each coordinate and replacing $i$ by i-i in the expression of the operator $\mathbb{D}_{\lambda,\nu}^{N,N}$. Specifically, from the expression of the map $\Phi_{\text{Sol}}^{+}$, $\mathbb{D}_{\lambda,\nu}^{N,-N}$ can be obtained by swapping the $s$-coordinate with $(-1)^{s}$ times the (2Ns)(2N-s)-coordinate and replacing $i$ by i-i in the operator $\mathbb{D}_{\lambda,\nu}^{N,N}$. This proves Theorem 1.3 for m=Nm=-N.
@@
Note that the results of this section can be applied when solving Problems A and B for $|m|>N$. In fact, it suffices to solve them when $m>N$ and then use Proposition 8.2 to obtain the solution for m<Nm<-N.
@@ 9.1 Renormalized Gegenbauer polynomials
$C_{\ell}^{\mu}(z)$ $=\sum_{k=0}^{[\frac{\ell}{2}]}(-1)^{k}\frac{\Gamma(\ell-k+\mu)}{\Gamma(\mu)k!(\ell-2k)!}(2z)^{\ell-2k}.$
@@ 9.2 Algebraic properties of renormalized Gegenbauer polynomials and imaginary Gegenbauer operators
Let $\mu\in{\mathbb{C}}$ and $\ell\in{\mathbb{N}}$ and suppose that $2\mu+\ell=0$. Then, for any $n,d\in{\mathbb{N}}$ we have:
@@
The proof is straightforward from the definition of $S^{\mu}_{\ell}$ and the condition $2\mu+\ell=0$. ∎